3x√8x - 4√x^3 + 3x√72x ;X=2.
x≠-y 3x/1/x+y (3x is over 1 which is over x+y =?)
1
Expert's answer
2012-08-21T09:30:02-0400
3x*sqrt(8x) + 4*sqrt(x^3) + 3x*sqrt(72x) = 6x*sqrt(2x) + 4x*sqrt(x) + 6*3x*sqrt(2x) = 24x*sqrt(2x) + 4x*sqrt(x);

If x=2 the value of expression is 24*2*2 - 8*sqrt(2) = 96 - 8*sqrt(2)